A ball with a mass of 275 g is dropped from rest, hits the floor and rebounds upward. If the ball hits the floor with a speed of 2.30 m/s and rebounds with a speed of 2.00 m/s, determine the following.

(a)magnitude of the change in the ball's momentum in kg · m/s (Let up be in the positive direction.)

kg · m/s

(b)change in the magnitude of the ball's momentum in kg · m/s (Let negative values indicate a decrease in magnitude.)

kg · m/s
Answer #1

Given,

The mass of a ball, m = 275 g = 0.275 Kg

The speeds, v1 = 2.30 m/s, and v2 = 2 m/s

(a)

The change in momentum, \Deltap = mv1 - mv2

\Deltap = m (v1 - v2)

= 0.275 ( 2.30 - 2)

= 0.0825 Kg. m/s

(b) The change in magnitude of momentum, \Deltap = | m (v1 - v2) |

\Deltap = | 0.275 ( 2.30 - (- 2)) |

= 1.1825 Kg m/s